# Mechanisms vs. Outcomes: Probing for Syntax Fails to Explain Performance on Targeted Syntactic Evaluations

### v1

@@ Abstract
Large Language Models (LLMs) exhibit a robust mastery of syntax when processing and generating text. While this suggests internalized understanding of hierarchical syntax and dependency relations, the precise mechanism by which they represent syntactic structure is an open area within interpretability research. Probing provides one way to identify the mechanism of syntax being linearly encoded in activations, however, no comprehensive study has yet established whether a model’s probing accuracy reliably predicts its downstream syntactic performance. Adopting a “mechanisms vs. outcomes” framework, we evaluate 32 open-weight transformer models and find that syntactic features extracted via probing fail to predict outcomes of targeted syntax evaluations across English linguistic phenomena. Our results highlight a substantial disconnect between latent syntactic representations found via probing and observable syntactic behaviors in downstream tasks.
@@ 1 Introduction
The remarkable ability of LLMs to process and generate text that respects a rich diversity of syntactic constraints strongly suggests that models have sophisticated syntactic knowledge. Yet our understanding of internal representations of these syntactic facts is lacking, which has motivated research into interpretability paradigms searching for syntactic mechanisms, such as mechanistic interpretability and causal abstractions Geiger et al. (2022); Murty et al. (2022). We study probing, a prominent approach in this domain which assumes that model activations are the mechanism through which syntactic knowledge is encoded and that this knowledge is linearly recoverable using small supervised models called probes (Conneau et al., 2018; Hewitt and Manning, 2019; Tenney et al., 2019, among others). However, no comprehensive study has evaluated whether probing accuracy is indicative of targeted syntactic outcomes.
@@
We deepen understanding of the relationship between syntactic mechanisms and outcomes by asking: Do probing accuracies, measured by tree attachment scores of syntactic probes, effectively predict a model’s downstream performance on targeted syntactic evaluations? We evaluate syntactic outcomes using BLiMP (Warstadt et al., 2020), a benchmark of minimal pairs spanning well-established English grammatical contrasts. Assigning higher probability to the acceptable sentence in an acceptable-unacceptable pair has been established as a desired linguistic outcome Marvin and Linzen (2018); Warstadt et al. (2020).
We train three syntax probes on 32 open-weight transformer models. We then fit Ordinary Least Squares (OLS) regressions modeling BLiMP minimal pairs accuracy as a function of probe attachment scores on the grammatically acceptable sentences (Fig 1). We demonstrate that across all levels of analysis—amalgamation of all BLiMP linguistic phenomena, individual phenomena, and targeted suites of challenging minimal pairs—syntax probing accuracy shows no clear predictive power for downstream outcomes. [^1] Code will be uploaded to GitHub soon.
Concretely, we find that probing performance—measured via directed and undirected unlabeled attachment scores (UAS/UUAS)—fails to predict model accuracy on BLiMP minimal pairs. At the full-dataset level, regression fits are negligible, with no statistically significant relationships across 32 models and three probes. At the per-phenomenon level, only anaphor agreement shows a significant correlation, but follow-up controls reveal that even a non-syntactic probe can produce similar effects, suggesting the result is likely spurious. In finer-grained evaluations targeting challenging syntactic paradigms, we test whether probes can recover critical edges—such as the subject-to-verb dependency in agreement examples—that are plausibly relevant for resolving minimal pairs. However, even when the probe identifies these edges correctly, model success on the corresponding minimal pair is only weakly aligned (40–60% match rate across models). Finally, at the sentence level, UUAS scores are largely overlapping between correctly and incorrectly evaluated minimal pairs in over 85% of BLiMP suites. Taken together, our results suggest that while syntactic probing is popular, it fails to robustly reveal a model’s latent syntactic knowledge—external targeted evaluations remain the gold standard.
@@ 2 Setup
Given an input sentence $S\triangleq(w_{1},\ldots,w_{N})$, a pre-trained transformer $f_{\theta}$ produces contextual vectors $h_{i}\in\mathbb{R}^{d{}}$ for each word $w_{i}$ at every layer. Our probes $g_{\phi}{}:\mathbb{R}^{d{}}\to\mathbb{R}^{k{}}$, $k{}\leq d{}$ are learned functions that project contextual hidden states linearly into a $k{}$-dimensional sub-space. Linear and bilinear probes have been shown to achieve high selectivity; they effectively capture linguistic properties of the probed representation (Hewitt and Liang, 2019).
@@ 2.1 Mechanisms
Proposed in Hewitt and Manning (2019), $g^{\text{struct}}_{\phi}$ is a linear transformation ($\phi=\{B^{\text{struct}}\in\mathbb{R}^{d{}\times k{}}\}$) that learns to encode the tree distance $d_{ij}$ between word pairs $w_{i}$ and $w_{j}$ in the probe’s projected space by minimizing
$g^{\text{struct}}_{\phi}(h)=B^{\text{struct}}h$
@@
The test metric is undirected unlabeled attachment score (UUAS): the fraction of gold tree edges included in the minimum spanning tree calculated from the probe’s predicted tree distances. Punctuation and the root relation are excluded.
Orthogonal Structural Probe
Limisiewicz and Mareček (2021) replace the $g^{\text{struct}}_{\phi}$ linear transformation with an orthogonal transformation and scaling vector $g^{\text{ortho}}_{\phi}$ where parameters of the probe $\phi=\{V\in\mathbb{R}^{d{}\times d},\bar{d}\in\mathbb{R}^{d}\}$. $V$ is orthogonal, and the authors maintain its orthogonality during training using Double Soft Orthogonality Regularization (DSO) Bansal et al. (2018) with regularization $\lambda_{o}=0.05$. The overall probe training objective is
$g^{\text{ortho}}_{\phi}(h)=\bar{d}\odot Vh$
@@
$+\lambda_{o}\text{DSO}(V).$ (2)
Head Word Probe
Inspired by Clark et al. (2019), $g^{\text{head}}_{\phi}$ ($\phi=\{B^{\text{head}}\in\mathbb{R}^{d{}\times k{}}\}$) formulates a classification task with cross-entropy loss: for each $w_{i}$, including punctuation, predict the head word in $\{\text{ROOT},w_{1},\ldots,w_{N}\}\setminus\{w_{i}\}$. Let $H$ be cross entropy loss and $\text{head}_{i}$ be the head word for $w_{i}$:
$g^{\text{head}}_{\phi}(h)=B^{\text{head}}h$
@@
$\min_{\phi}\sum_{S}\frac{1}{|S|}\sum_{i}H(\{\hat{d}_{ij}\mid i\neq j\},\text{head}_{i}).$ (3)
Since the edge prediction is directed, the test metric is unlabeled attachment score (UAS).
@@
Importantly, training syntactic probes alone is insufficient to draw conclusions about their ability to predict syntactic evaluation performance because the quality of model hidden states is a confounding variable affecting both probe and syntactic evaluation performance. To isolate this confounder, we propose a simple control probing task where performance depends only on hidden state quality and not on latent syntactic structure. Concretely, for each sentence, we form all word pairs where the treebank-specific part-of-speech (XPOS) of the two words is the same and train a probe to recover the $L_{2}$ distance between the two words’ GloVe vectors Pennington et al. (2014). Let $v_{i}$ denote the GloVe vector of $w_{i}$. The training objective for the control probe is
@@
$\min_{\phi}\sum_{S}\frac{1}{|S|}\sum_{\mathclap{\begin{subarray}{c}(i,j)\\ \text{POS}(w_{i})=\text{POS}(w_{j})\end{subarray}}}\;\text{Huber}(\tilde{d}_{ij},\|v_{i}-v_{j}\|_{2}).$ (4)
@@
This probe is trained to recover non-syntactic information as modeled by an uncontextualized word embedding model, GLoVe, whose vector geometries have been shown to encode linguistic information like lexical semantics (Brown et al., 2023). We further ensure that our control probe is not indirectly capturing syntactic part-of-speech information by only training the probe to recover within word-pair distances where the words share the same XPOS. Crucially, any predictive success of the control probe on syntactic evaluations warrants a more cautious interpretation of the results of syntactic probes, as it suggests that the latter’s explanatory power may not be attributable solely to the recoverability of syntactic representations.
@@ 2.2 Outcomes
Our benchmark for linguistic knowledge outcomes is BLiMP accuracy. BLiMP has 67 template-generated sub-datasets called paradigms grouped into 13 linguistic phenomena across English morphology, syntax, and semantics Warstadt et al. (2020). [^2] The two semantic phenomena, quantifiers and NPI licensing, still require knowledge of hierarchical syntactic structure, e.g., to determine scope for NPI licensing (Ladusaw, 1979). Each paradigm has 1000 acceptable-unacceptable minimal pairs $(S_{acc},S_{unacc})$. Accuracy on a set of pairs $\mathcal{D}$ is computed as
@@
For decoder models, $\log P_{f_{\theta}}(S)$ is the sum of token level log-probabilities. For encoder and encoder-decoder models, we instead use pseudo-log-likelihood scores (PLL; Salazar et al., 2020).
@@ 3 Experimental Setup
We train probes on Stanza Qi et al. (2020) Universal Dependencies Zeman et al. (2019) dependency parses of the Penn Treebank (PTB; Marcus et al., 1993) using its standard train and validation splits. We report training parameters and compute in Appendix B. $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$ have fixed probe dimension $k{}=256$, while $g^{\text{ortho}}_{\phi}$ has $k=d$ for each model to ensure orthogonality is feasible. Let $L$ be the number of layers in LM $f_{\theta}$. To reduce the influence of a word’s semantic properties, we subtract word embeddings from each layer’s contextual states
@@
Training produces $L$ different probes for each model. For each of $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$, we select the probe with the best PTB test metric. For $g^{\text{ortho}}_{\phi}$, given limited compute resources and the fact that it is derived from $g^{\text{struct}}_{\phi}$, we train a single probe per model on the layer with the best $g^{\text{struct}}_{\phi}$ test metric. For $g^{\text{ctrl}}_{\phi}$ we train at most two probes per model corresponding to the best layers for $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$, which may match.
@@
We construct a simple evaluation dataset to test how well our control probe abstracts away contextual linguistic information to verify its design. For 500 random words chosen from SimLex999 (Hill et al., 2014), we prompt GPT-4o OpenAI et al. (2024) to generate 10 sentences that contain the word in substantially different contexts. Refer to Appendix C for the prompt and sample sentences.
@@
Our objective is model-level mechanisms vs. outcomes correlation because we want to study how effectively the method of syntactic probing reveals an arbitrary model’s latent syntactic knowledge. We use the PTB test set scores to choose the best probe from each probe family and apply it to every $S_{acc}$ from BLiMP to compute UUAS/UAS/$\rho_{s}$ as predictor variables (mechanisms). $S_{acc}$ and $S_{unacc}$ feed into minimal pair evaluation as the response variable (outcomes).
We fit simple (eqn 6) and multiple (eqn 7) OLS regressions.
@@
$y$ is minimal pairs accuracy (eqn 5). We run simple regression with each probe individually: $x_{1}$ is either BLiMP UUAS, UAS, or Spearman correlation $\rho_{s}$. For multiple regression, $x_{1}$ is UUAS or UAS and $x_{2}$ is $\rho_{s}$. Comparing $\beta_{1}$ and its $p$-value for the syntax probe when it is the sole predictor versus when paired with the control reveals the extent to which it predicts grammaticality outcomes. Since the two regression models are nested, we also include log-likelihood ratio tests (LRT) for robust statistical comparisons.
We experiment at three BLiMP granularities: (1) full dataset (average results of all test suites), (2) each of the 13 phenomena separately with Holm-Bonferroni correction, (3) the two test suites in each of the subject-verb agreement and filler-gap phenomena that scored lowest in Warstadt et al. (2020). For the individual test suite experiments, rather than tree UUAS/UAS, we determine if the induced syntactic tree contains a critical edge corresponding to the phenomenon targeted by the minimal pairs. Appendix D discusses critical edge criteria. We hypothesize that the binary outcome of the probe’s ability to extract the critical edge on $S_{acc}$ is linked to the minimal pair binary outcome.

### v2

@@ Abstract
Large Language Models (LLMs) exhibit a robust mastery of syntax when processing and generating text. While this suggests internalized understanding of hierarchical syntax and dependency relations, the precise mechanism by which they represent syntactic structure is an open area within interpretability research. Probing provides one way to identify syntactic mechanisms linearly encoded in activations; however, no comprehensive study has yet established whether a model’s probing accuracy reliably predicts its downstream syntactic performance. Adopting a “mechanisms vs. outcomes” framework, we evaluate 32 open-weight transformer models and find that syntactic features extracted via probing fail to predict outcomes of targeted syntax evaluations across English linguistic phenomena. Our results highlight a substantial disconnect between latent syntactic representations found via probing and observable syntactic behaviors in downstream tasks.
@@ 1 Introduction
The remarkable ability of LLMs to process and generate text that respects a rich diversity of syntactic constraints strongly suggests that models have sophisticated syntactic knowledge. Yet our understanding of internal representations of these syntactic facts is lacking, which has motivated research into interpretability paradigms searching for syntactic mechanisms, such as mechanistic interpretability and causal abstractions geiger2022; murty2023intrinsic. We study probing, a prominent approach in this domain which assumes that model activations are the mechanism through which syntactic knowledge is encoded and that this knowledge is linearly recoverable using small supervised models called probes (conneau-etal-2018-cram; hewitt-manning-2019-structural; tenney-etal-2019-bert, among others). However, no comprehensive study has evaluated whether probing accuracy is indicative of targeted syntactic outcomes.
@@
We deepen understanding of the relationship between syntactic mechanisms and outcomes by asking: Do probing accuracies, measured by dependency tree attachment scores of syntactic probes, effectively predict a model’s downstream performance on targeted syntactic evaluations? We evaluate syntactic outcomes using BLiMP (warstadt-etal-2020-blimp-benchmark), a benchmark of minimal pairs spanning well-established English grammatical contrasts. Assigning higher probability to the acceptable sentence in an acceptable-unacceptable pair has been established as a desired linguistic outcome marvin-linzen-2018-targeted; warstadt-etal-2020-blimp-benchmark.
We train three syntax probes on 32 open-weight transformer models. We then fit Ordinary Least Squares (OLS) regressions modeling BLiMP minimal pairs accuracy as a function of probe attachment scores on the grammatically acceptable sentences (Fig 1). We demonstrate that across all levels of analysis—amalgamation of all BLiMP linguistic phenomena, individual phenomena, and targeted suites of difficult minimal pairs—syntax probing accuracy shows no clear predictive power for downstream outcomes. [^1] Code is at https://github.com/agananth/SyntaxMechanismsOutcomes
Concretely, we find that probing performance—measured via directed and undirected unlabeled attachment scores (UAS/UUAS)—fails to predict model accuracy on BLiMP minimal pairs. Across 32 models, none of the three probes yields a statistically significant regression fit on the full overall dataset. At the per-phenomenon level, only anaphor agreement shows a significant correlation, but the control experiment reveals that even a non-syntactic probe can produce a similar fit, suggesting the result is likely spurious. In finer-grained evaluations targeting difficult syntactic paradigms, we test whether probes can recover critical tree edges that are plausibly relevant for resolving minimal pairs, such as the subject-verb edge in subject–verb agreement pairs. However, probe recovery of these edges only weakly aligns with minimal pair outcome (40–60% match rate across models). Finally, within each BLiMP paradigm we test whether UUAS score distributions differ between correct and incorrect minimal pair outcomes. In over 85% of paradigms, the distributions overlap substantially for most models.
Taken together, our results suggest that while syntactic probing is a dominant interpretability paradigm, it fails to robustly reveal a model’s latent syntactic knowledge. We establish a disconnect between the syntactic information revealed through conventional probing for dependency trees and the latent syntactic knowledge that far more often than not assigns grammatical text a higher likelihood than ungrammatical text. Our results argue for the use of external targeted evaluations as the gold standard for establishing model syntactic competence, and support the continuing development of BLiMP-style resources beyond English (jumelet2025multiblimp10massivelymultilingual; başar2025turblimpturkishbenchmarklinguistic; taktasheva2024rublimprussianbenchmarklinguistic, among others).
@@ 2 Setup
Given an input sentence $S\triangleq(w_{1},\ldots,w_{N})$, a pre-trained transformer $f_{\theta}$ produces contextual vectors $h_{i}\in\mathbb{R}^{n{}}$ for each word $w_{i}$ at every layer. Our probes, denoted $g_{\phi}:\mathbb{R}^{n{}}\to\mathbb{R}^{k{}}$ with parameters $\phi$, are learned linear functions that project contextual hidden states into a kk{}-dimensional subspace, where $k{}\leq n{}$. Linear and bilinear probes have been shown to achieve high selectivity; they effectively capture linguistic properties of the probed representation (hewitt2019designinginterpretingprobescontrol).
@@ 2.1 Mechanisms
Proposed in hewitt-manning-2019-structural, the structural distance probe $g^{\text{struct}}_{\phi}$ is a linear transformation $g^{\text{struct}}_{\phi}(h)=B^{\text{struct}}h$ with parameters $\phi=\{B^{\text{struct}}\in\mathbb{R}^{k{}\times n{}}\}$ that learns to encode the dependency tree distance $d_{ij}$ between word pair $w_{i}$ and $w_{j}$ in the probe’s projected space by minimizing
@@
The evaluation metric is undirected unlabeled attachment score (UUAS): the fraction of gold tree edges $E_{\text{gold}}=\{\{w_{i},w_{j}\}:(w_{i},w_{j})\in\mathcal{G}\}$ included in the minimum spanning tree $E_{\text{pred}}=\{\{w_{i},w_{j}\}:(w_{i},w_{j})\in\hat{\mathcal{G}}\}$ calculated from the probe’s predicted tree distances. Punctuation and the root relation are excluded.
$\text{UUAS}=(|E_{\text{pred}}\cap E_{\text{gold}}|)/|E_{\text{gold}}|.$
Orthogonal Structural Probe.
Limisiewicz2021orth replace the original structural probe linear transformation with an orthogonal transformation and scaling vector. The orthogonal probe $g^{\text{ortho}}_{\phi}$ is the transformation $g^{\text{ortho}}_{\phi}(h)=\bar{d}\odot Vh$ with parameters $\phi=\{V\in\mathbb{R}^{n{}\times n},\bar{d}\in\mathbb{R}^{n}\}$ where $V$ is orthogonal. The authors maintain orthogonality during training using Double Soft Orthogonality Regularization (DSO; bansal2018gainorthogonalityregularizationstraining) with regularization $\lambda_{o}=0.05$. The overall probe training objective is
@@
[…]
Limisiewicz2021orth report $g^{\text{ortho}}_{\phi}$ performs on par with $g^{\text{struct}}_{\phi}$ at predicting dependency trees while being less prone to memorizing training trees. Moreover, the scaling vector $\bar{d}$ enables interpretation of the relative importance of each dimension in $V$. The evaluation metric is UUAS, as with $g^{\text{struct}}_{\phi}$.
Head Word Probe.
Inspired by clark2019bert, we define the head word probe $g^{\text{head}}_{\phi}$ as a linear transformation $g^{\text{head}}_{\phi}(h)=B^{\text{head}}h$ with parameters $\phi=\{B^{\text{head}}\in\mathbb{R}^{k{}\times n{}}\}$ trained using cross-entropy loss. For each dependent $w_{i}$ (including punctuation), the probe predicts its head in $\{\text{ROOT},w_{1},\ldots,w_{N}\}\setminus\{w_{i}\}$. Let $H$ be cross entropy loss and $\text{head}(i)$ be the head of $w_{i}$:
@@
$\min_{\phi}\sum_{S}\frac{1}{|S|}\sum_{i}H(\{\hat{d}_{ij}\mid i\neq j\},\text{head}(i)).$ (3)
Since the edge prediction is directed, the evaluation metric is unlabeled attachment score (UAS). Here $E_{\text{gold}}=\{(i,\mathrm{head}(i)):1\leq i\leq N\}$ and $E_{\text{pred}}=\{(i,\hat{\mathrm{head}}(i)):1\leq i\leq N\}$ such that
$\text{UAS}=(|E_{\text{pred}}\cap E_{\text{gold}}|)/|E_{\text{gold}}|.$
@@
Importantly, the performance of syntax probes and downstream syntactic performance may partly reflect non-syntactic factors such as the quality of lexical semantic representations in a model. The overall quality of model hidden states is thus a confounding variable affecting both probe and syntactic evaluation performance. To isolate this confounder, we propose a simple control probing task where performance depends only on hidden state quality and not latent syntactic structure. Concretely, for each sentence, we form all word pairs where the treebank-specific part-of-speech (XPOS) of the two words is the same and train a probe $g^{\text{ctrl}}_{\phi}$ to recover the $L_{2}$ distance between the two words’ GloVe vectors pennington-etal-2014-glove. $g^{\text{ctrl}}_{\phi}$ is a linear transformation $g^{\text{ctrl}}_{\phi}(h)=B^{\text{control}}h$ with parameters $\phi=\{B^{\text{control}}\in\mathbb{R}^{k{}\times n{}}\}$. Let $v_{i}$ denote the GloVe vector of $w_{i}$. The training objective for the control probe is
@@
$\min_{\phi}\sum_{S}\frac{1}{|S|}\sum_{\mathclap{\begin{subarray}{c}(i,j)\\ \text{XPOS}(w_{i})=\text{XPOS}(w_{j})\end{subarray}}}\;\text{Huber}(\tilde{d}_{ij},\|v_{i}-v_{j}\|_{2}).$ (4)
@@
This probe is trained to recover non-syntactic information as modeled by an uncontextualized word embedding model, GloVe, whose vector geometries have been shown to encode linguistic information like lexical semantics (https://doi.org/10.1111/cogs.13291). We further ensure that our control probe is not indirectly capturing syntactic part-of-speech information by only training the probe to recover within word-pair distances where the words share the same XPOS. Crucially, any predictive success of the control probe on syntactic evaluations warrants a more cautious interpretation of the results of syntactic probes, as it suggests that the latter’s explanatory power may not be attributable solely to the recoverability of syntactic representations.
@@ 2.2 Outcomes
Our benchmark for linguistic knowledge outcomes is BLiMP accuracy. BLiMP has 67 template-generated sub-datasets called paradigms grouped into 13 linguistic phenomena across English morphology, syntax, and semantics warstadt-etal-2020-blimp-benchmark. [^2] The two semantic phenomena, quantifiers and NPI licensing, still require knowledge of hierarchical syntactic structure, e.g., to determine scope for NPI licensing (Ladusaw1979). Each paradigm has 1000 acceptable-unacceptable minimal pairs $(S_{acc},S_{unacc})$ that exhibit a single grammatical contrast. The example pair from Fig 1 is from the distractor agreement relative clause paradigm within subject–verb agreement. The verb agreement contrast point in the pair between the acceptable (top) and unacceptable (bottom) sentences is shown below in bold:
[…]
Model $f_{\theta}$ accuracy on a set of pairs $\mathcal{D}$ is computed as
@@
For decoder models, $\log P_{f_{\theta}}(S)$ is the sum of token level log-probabilities. For encoder and encoder-decoder models, we instead use pseudo-log-likelihood scores (PLL; Salazar_2020).
@@ 3 Experimental Setup
We train each of our four probe “families” on Stanza qi-etal-2020-stanza Universal Dependencies de-marneffe-etal-2021-universal dependency parses of the Penn Treebank (PTB; marcus-etal-1993-building) using its standard train and validation splits. We report training parameters and compute in Appendix B. $g^{\text{struct}}_{\phi}$, $g^{\text{head}}_{\phi}$, and $g^{\text{ctrl}}_{\phi}$ have fixed probe dimension k=256k{}=256, while $g^{\text{ortho}}_{\phi}$ has $k=n$ for each model to ensure orthogonality is feasible. Let $L$ be the number of layers in model $f_{\theta}$. To reduce the influence of a word’s semantic properties, we subtract word embeddings from each layer’s contextual states
@@
For each of $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$, we train $L$ probes per model—one per layer—and select the probe with the best PTB test metric. For $g^{\text{ortho}}_{\phi}$, given limited compute resources and the fact that it is derived from $g^{\text{struct}}_{\phi}$, we train a single probe per model on the layer with the best $g^{\text{struct}}_{\phi}$ PTB test metric. Since the best layer can differ between $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$, and the objective of $g^{\text{ctrl}}_{\phi}$ is to control for hidden state quality, we train a separate $g^{\text{ctrl}}_{\phi}$ instance for each on the corresponding best layer. Control results are reported with the associated probe.
@@
We construct a simple evaluation dataset to test how well our control probe abstracts away contextual linguistic information to verify its design. For 500 random words chosen from SimLex999 (hill2014simlex999evaluatingsemanticmodels), we prompt GPT-4o openai2024gpt4ocard to generate 10 sentences that contain the word in substantially different contexts. Refer to Appendix C for the prompt and sample sentences.
@@
Our objective is model-level mechanisms vs. outcomes regression because we want to study how effectively the method of syntactic probing predicts an arbitrary model’s latent syntactic knowledge. We compute regressions separately at two BLiMP granularities: (1) overall (averaging predictor and response values across all paradigms in the full dataset), and (2) each of the 13 phenomena separately with Holm-Bonferroni correction. Holm-Bonferroni is an effective method for adjusting significance tests to account for multiple comparisons (13 in our case).
For each linear regression at a given granularity, the predictor variables are probe scores: UUAS ($g^{\text{struct}}_{\phi}$, $g^{\text{ortho}}_{\phi}$), UAS ($g^{\text{head}}_{\phi}$), or Spearman correlation $\rho_{s}$ ($g^{\text{ctrl}}_{\phi}$). We run regressions separately for each probe family, with one data point per model. Fig 1 illustrates the approach: the selected probe for each model is applied to every $S_{acc}$ sentence at the given granularity to compute an average probe score (mechanism), and each $S_{acc}$ and $S_{unacc}$ feed into minimal pair evaluation (outcome).
We fit simple (eqn 6) and multiple (eqn 7) OLS regressions for each probe family at each granularity:
@@
where $y$ is minimal pairs accuracy (eqn 5) and $x_{1}$ is the probe score. For multiple regression, $x_{1}$ is UUAS or UAS and $x_{2}$ is $\rho_{s}$. Comparing $\beta_{1}$ and its $p$-value for the syntax probe when it is the sole predictor versus when paired with the control reveals the extent to which it predicts grammaticality outcomes. Since the two regression models are nested, we also include log-likelihood ratio tests (LRT) for robust statistical comparisons.
Critical Edge Study in Difficult Paradigms.
Drilling down to the paradigm level, we select the two paradigms in each of the subject–verb agreement (distractor agreement relational noun, distractor agreement relative clause) and filler–gap (wh. vs. that with gap, wh. vs. that with gap long distance) phenomena that scored lowest in model accuracy in warstadt-etal-2020-blimp-benchmark. Rather than tree UUAS/UAS, we determine if the induced syntactic tree for each $S_{acc}$ contains a critical edge corresponding to the phenomenon targeted by the minimal pairs—an nsubj edge for subject–verb agreement, and an obj or obl edge for filler–gap. Appendix D discusses critical edge criteria in-depth. We hypothesize that the binary outcome of the probe’s ability to extract the critical edge on $S_{acc}$ is linked to the minimal pair binary outcome.